OpenStudy (mackenzie2013):

A man riding a motorbike travels a distance of 285 miles in 6 hours. What is the motorbike's speed in feet per minute? 4810 ft/min 588 ft/min 855 ft/min 4180 ft/min

OpenStudy (anonymous):

\[\frac{285* 5280}{6* 60}=\frac{4180 \text{ ft}}{\min } \]